# Interacting tachyon with varying mass dark matter

### v1

@@ Abstract
Tachyon; Interaction; Varying mass Dark matter, Phase space analysis, Center manifold theory, State finder Stability
pacs:
95.36.+x, 95.35.+d, 98.80.-k, 98.80.Cq.
@@ I Introduction
Inflation provides a technique for production of density perturbations which is required to mould the evolution of the universe. In the simple case of inflation, the universe is dominated with a scalar field and potential energy dominates over the kinetic term, followed by a reheating period [12, 13]. But there is no well accepted exposition to integrate inflationary scenario, the scalar field that drives inflation, with one of the known fields of particle physics. It is also salient to emerge inflation potential naturally from underlying fundamental theory. In this regard, tachyon fields analogous to unstable D-branes could be accountable for inflation in early time [14, 15]. In order to address such issues, many theoretical models of DE have been raised in the theoretical background among them interacting cosmological models where DE interacts with DM have gained a remarkable attention with a successive number of observational data [6]. Infact, latest observations indicate that there could be an non-vanishing interaction within $1\sigma$ confidence region in dark sectors [16, 17, 18]. Due to unknown nature of two dark components there is no standard form of interaction term even there is no any guiding principle to choose an interaction between them. Therefore, one can adopt the interaction purely on the basis of phenomenology. In fact, an appropriate interaction can provide a good mechanism to alleviate the coincidence problem. The models of DE interacting with DM have extensively been studied in the literature [25, 24, 26, 27, 19, 20, 21, 22, 23, 28, 29, 30, 31, 32, 33]. As nothing much is known about the physical nature of DE, probe is still on to hunt for an appropriate candidate for DE.
@@
On the other hand, interacting DE with varying mass DM particles can also provide a proper way to alleviate coincidence problem. The model of varying mass DM particles is based on the mechanism that decaying of DM reproduces the scalar field as DE. Here, mass of the dark matter is dependent on cosmic time ‘$t$’ via a scalar field ‘$\phi$’. It should be mentioned that this mechanism can be treated equivalent to an interaction included in the dark sectors, see for example [34, 35]. However, the model of varying mass has some basic differences from interacting models. The mass of dark matter can be exponential or power-law functions of scalar field or any other general form. Interacting quintessence studied recently with varying mass dark matter in the Lyra’s manifold in context of dynamical analysis, where mass of dark matter is assumed to be vary a exponential function of scalar field [36]. Interacting phantom is also studied where exponential and power-law of scalar field as well as exponential or power-law mass dependence is undertaken in the perspective of dynamical system [37]. Consequently, a center manifold theory is employed to study the interacting phantom in association with varying mass dark matter [38]. A plethora of interacting quintessence models have been put forward over the years in the literature [39, 40, 41] where the dark matter mass depends on exponential or power-law of scalar field $\phi$, and these provided the possible solutions to coincidence problem. In this context, we would like to mention that the tachyon field can be made as a satisfactory candidate for the high energy inflation [42] and at the same time as an origin of DE depending on the form of the tachyon potential [43].
@@
The paper is organized as follows. In the next Sec. II, we first discuss about the cosmological model considered here and then discuss about the formation of autonomous system and the relevant parameters. Phase space analysis has also been presented in Sec. III. In addition, in Sec. IV, we examine the behavior of diagnostic pair parameters. In Sec. V we depicts the evolution of different cosmological quantities for our model and compare it with the observational data. Finally, our main findings and conclusions are shortly discussed in Sec. VI. Throughout the paper, we use natural units in which $\kappa^{2}=8\pi G=c=1$.
@@ II.1 The model of interacting varying mass tachyonic DE
As is well-known, the universe is supposed to be homogeneous and isotropic at large scale and the spatially flat Friedmann-Lemaître-Robertson-Walker (FLRW) metric describes it very well. Here, we start with the FLRW metric given by the line element
@@
where $\rho_{m}$ denotes the energy density of the DM. The energy density and the thermodynamic pressure for the tachyonic DE are denoted by $\rho_{\phi}$ and $p_{\phi}$ and are defined as follows:
@@
$Q=\frac{Q_{m}^{{}^{\prime}}(\phi)}{Q_{m}(\phi)}\dot{\phi}\rho_{m}=\beta(\phi)\dot{\phi}\rho_{m}$ (17)
where $\beta(\phi)=\frac{\partial}{\partial\phi}ln(Q_{m}(\phi))$. It should be noted that Honorez et al. [35] studied a coupled quintessence model in which the interaction with the dark matter sector is a function of the quintessence potential. They have also showed that such type of interaction (17) can arise from a field dependent mass term for the DM field. This kind of coupling (i.e., $\beta(\phi)=$ constant) can be found in the literature where non-minimally coupled Brans-Dicke theory containing a self-interacting potential, can automatically give rise to an interaction between the Brans-Dicke scalar field and the normal matter by applying a suitable conformal transformation (see [31] and the references therein). In our case, it is also clearly seen that the tachyonic scalar field and the DM do not evolve independently but interact with each other via an interaction term $Q$ in the energy conservation Eqs. (15) and (16). It is important to note here that if we concentrate on the other interacting DE-DM models, the form of coupling chosen is ad-hoc and the source of such coupling is not known (see e.g. [19]). But, in our case, unlike other models, the interaction term is not an input but obtained its form from the field equations by considering the variable mass particle scenarios as already referred earlier. This makes our work very interesting and deserves further study in the present context.
Moreover, one can also express the conservation equations (15) and (16) in non-interacting form with effective EoS parameters for DM and DE as follows:
@@
The dynamics of cosmological evolution of the universe in this scenario of VAMP, we will perform the critical points analysis in the next section.
@@ II.2.1 Autonomous system with exponential mass function and exponential potential
In the model of exponential variable mass particles of DM, we consider the mass of DM is function of scalar field $\phi$ as $Q_{m}(\phi)=Q_{m0}~{}\mbox{exp}\{-\alpha H_{0}\phi\}$ and the potential of the scalar field $\phi$ as $V(\phi)=V_{0}~{}exp\{\beta H_{0}\phi\}$, where $Q_{m0}$ and $V_{0}$ are constant and $\alpha$ and $\beta$ are constant parameters. Now with the exponential potential and exponential mass function of DM, the 5D autonomous system (22) reduces to the following 3D autonomous system of ordinary differential equations:
@@
It is clear that the autonomous system (23) has singularities at $y=0$ and $z=1$. In this situation we can not study the properties of solutions lying on $y=0$ and $z=1$ plane. In order to remove the singularities, we multiply the right hand sides of the system (23) by $y^{2}(1-z)$. This operation allows us to analyze the solutions on $y=0$ and $z=1$ plane without changing the qualitative dynamical features of the system in the other regions of the phase space for $y^{2}(1-z)>0$. After applying this technique, we obtain
@@
The effective equation of state parameter for tachyon field (DE) read as
@@
From the above, one can achieve the condition for acceleration of the universe when: $q<0$, i.e., when $\omega_{eff}<-\frac{1}{3}$ and for deceleration, one follows the condition: $q>0$, i.e., $\omega_{eff}>-\frac{1}{3}.$ Friedmann equation (2) will give the constraint equation of $\Omega_{m}$ explicitly depending on the variables $x$ and $y$ as
@@
III Phase space analysis of autonomous system (24):
The critical points for the system (24) are the following
@@ III Phase space analysis of autonomous system (24):
For autonomous system (24), Critical points and their corresponding physical parameters are shown in the table (1).
Table 1: The Critical points and the corresponding physical parameters for the system (24) are presented
@@
Table 2: The eigenvalues of the linearized system (24)
@@
The Table 2 shows the eigenvalues of the Jacobian matrix (evaluated at the critical points) for the autonomous system (24) and we note that all the critical critical points are nonhyperbolic in nature. So we can not analyze them by linear stability theory. We need to use center manifold theory for some cases. But center manifold theory can be used only for the critical points $B$ and $C$ as rest of the critical points do not have any eigenvalues with nonzero real part. To analyze $A,D,E,F,G$ we use numerical procedure.
@@
It is to be noted that the origin is a saddle on the $xy$ plane (see Fig. (1)) and the vector field near $x$-axis ($A$) is unstable on both $xy$ and $xz$ plane (see Fig. (1)). The vector field near the critical point $D$ on the $xy$-plane behaves as unstable node. The flow is repelling along $x$-axis and $y$-axis on the $xy$ plane. On the other hand, the line $(1,0,z_{c})$ is an attractor for $z_{c}>0$ and repeller for $z_{c}<0$ on the $xz$ plane. But the line $(1,0,z_{c})$ is a repeller on the $yz$ plane. In figure (3) it is to be noted that the vector field on the $xy$ plane near the critical point $E$ is unstable in nature. On the $xz$ plane the vector field is attracting for $z_{c}<0$ and repelling for $z_{c}>0$.
@@
$\bullet$ The set of critical points $B$ and $C$ exist for all $\beta\in\mathbb{R}$ and $\alpha$. These solutions are completely dominated by scalar field $(\Omega_{\phi}=1)$ which behaves as cosmological constant ($\omega^{\phi}_{eff}=-1$). The cosmic evolution near the points characterizes the de-Sitter expansion ($q=-1$) of the universe. These critical points are isolated nonhyperbolic in nature and the Jacobian matrix at $B$ or $C$ contains only one vanishing eigenvalue. We can employ center manifold theory to find the characteristic of the vector field near these critical points as the following:
@@
The equation (24) reduces to
$\frac{dx}{dN}$ $=-3x-\beta z-6xy+3xz-2(\alpha+\beta)zy+\text{higher degree terms}.$ (33)
$\frac{dy}{dN}$ $=-3y-\frac{3}{2}y^{2}+\frac{\beta}{2}xz+3zy+\text{higher degree terms}.$ (34)
$\frac{dz}{dN}$ $=-3zy-\frac{15}{2}zy^{2}+3z^{2}y+\frac{3}{4}zx^{2}+\text{higher degree terms}.$ (35)
The Jacobian matrix of the system (33-35) at the origin is given by
$J(0,0)=\begin{bmatrix}-3&0&\beta\\ 0&-3&0\\ 0&0&0\end{bmatrix}.$
To find center manifold at the origin, we need to diagonalize the above Jacobian matrix. For that we need to find $P$ such that $P^{-1}JP$ is a diagonal matrix. So $P$ is given by
$P=\begin{bmatrix}1&0&-1\\ 1&1&0\\ 0&0&\frac{3}{\beta}\end{bmatrix}$
If we consider J(0,0) as a linear transformation from $\mathbb{R}^{3}$ to $\mathbb{R}^{3}$ and the basis of $\mathbb{R}^{3}$ changes by $P$, then the co- ordinate vector (variables) $(x,y,z)$ transform to $(X,Y,Z)$ by the following rule
$\begin{bmatrix}x\\ y\\ z\end{bmatrix}=\begin{bmatrix}X-Z\\ X+Y\\ \frac{3Z}{\beta}\end{bmatrix}$ (42)
@@
$\begin{bmatrix}X\\ Y\\ Z\end{bmatrix}=\begin{bmatrix}x+\frac{\beta z}{3}\\ -x+y-\frac{\beta z}{3}\\ \frac{\beta z}{3}\end{bmatrix}$ (49)
Now for the variables $(X,Y,Z)$ the system (33-35) takes the following form:
$\frac{dX}{dN}$ $=-3X-6X^{2}-\frac{9}{\beta}Z^{2}-6XY-3\left(1-\frac{3-2\alpha}{\beta}\right)ZX-3\left(\frac{\beta+2\alpha}{\beta}\right)ZY+\text{higher degree terms}.$ (50)
$\frac{dY}{dN}$ $=-3Y+\left(\frac{9}{2}+\beta\right)X^{2}-\frac{3}{2}Y^{2}-\left(\frac{3}{2}-\frac{9}{\beta}\right)Z^{2}+\left(3+\beta\right)XY+\left(\frac{6\alpha}{\beta}+\frac{3}{2}-\beta\right)ZX$ (51)
$-\left(\beta-\frac{9+6\alpha}{\beta}\right)ZY+\text{higher degreeterms}.$ (52)
$\frac{dZ}{dN}$ $=-3Z\left(X+Y\right)+\text{higher degree terms}.$ (53)
According to center manifold theory we have the following
$X$ $=$ $-\frac{3}{\beta}Z^{2}+\text{higher degree terms}.$ (54)
$Y$ $=$ $(\frac{3}{\beta}-\frac{1}{2})Z^{2}+\text{higher degree terms}.$ (55)
Then the flow along the center manifold at $B$ can be written as
$\frac{dZ}{dN}=\frac{3}{2}Z^{3}+\text{higher degree terms}.$ (57)

### v2

@@ Abstract
Tachyon; interaction; varying mass dark matter, phase space analysis, Center manifold theory, Statefinder stability.
@@ I Introduction
Inflation provides a technique for production of density perturbations which is required to mould the evolution of the universe. In the simple case of inflation, the universe is dominated with a scalar field and potential energy dominates over the kinetic term, followed by a reheating period [12, 13]. But there is no well accepted exposition to integrate inflationary scenario, the scalar field that drives inflation, with one of the known fields of particle physics. It is also salient to emerge inflation potential naturally from underlying fundamental theory. In this regard, tachyon fields analogous to unstable D-branes could be accountable for inflation in early time [14, 15]. It has been first shown in Ref.[16] that the mechanism of decaying of non-BPS D4-brane into a stable D3 brane will give rise to a tachyon field via tachyon condensation, and that can act as a cosmological inflating field. After that, Sen in Ref.[17, 18, 19] showed that rolling tachyon field can arise naturally from unstable D-branes in bosonic and superstring theories. Further, he showed in Ref.[19] that the effective field of such tachyon may be designed as Dirac-Born-Infeld (DBI) scalar field theory. Furthermore, investigations of DBI field have been carried out in Refs.[20, 21] in the framework of FLRW cosmology to give inflating field behaviour.
It should be mentioned that even in bosonic string theory, the quantization of the Polyakov action (conformal transformation of Nambu-Goto action) provides a tachyonic like field giving locally unstable nature due to reverse sign of kinetic term, however, which soon decays through spontaneous symmetry breaking [22, 23]. In the present work, we have undertaken DBI field as effective field theory D-branes, so, there are no more instability problems as all the instable modes have been integrated out as shown in references [18, 19].
In order to address such issues, many theoretical models of DE have been raised in the theoretical background among them interacting cosmological models where DE interacts with DM have gained a remarkable attention with a successive number of observational data [6]. Infact, latest observations indicate that there could be an non-vanishing interaction within $1\sigma$ confidence region in dark sectors [24, 25, 26]. It should also be mentioned that the authors in Ref. [27] have tested a phenomenological interacting model with Planck data. After that another interacting model in Ref.[28] is tested by using SNeIa, BAO measurements, Hubble parameter data and CMB observations. Further, an interaction term with linear combination of energy densities of dark matter and dark energy was investigated in Ref. [29] with the SNeIa data. It was shown that for a small value of the coupling of interaction, the model tends to $\Lambda$CDM.
Due to unknown nature of two dark components there is no standard form of interaction term even there is no any guiding principle to choose an interaction between them. Therefore, one can adopt the interaction purely on the basis of phenomenology. In fact, an appropriate interaction can provide a good mechanism to alleviate the coincidence problem. The models of DE interacting with DM have extensively been studied in the literature such as an interacting quintessence is investigated in Ref.[30] where the interaction terms are taken in terms of energy densities and other covariant quantities. A detailed analysis of interacting quintessence is performed at the perturbative level (see in Ref. [31]). Linear and non-linear interactions have been investigated in context of phantom cosmology [32]. A class of linear and non-linear interactions have been investigated in Refs. [33, 34] where the interaction terms depend on energy densities of dark energy, dark matter and their derivatives. Interacting Chaplygin gas with different linear and non-linear sign changeable interactions are studied in Ref.[35]. For a detailed review on linear and non-linear interactions, one can follow the review paper [36]. Further, cosmological models interacting scalar field are investigated in modified gravity in Refs. [37, 38]. In Ref. [39], the authors have studied the interaction between dark energy-dark matter in the framework of irreversible thermodynamics of open system with particle creation. Interacting model of dark energy in Brans Dicke theory was investigated in Ref. [40] and interacting Barrow holographic dark energy is investigated in Ref. [41]. Recent studies (see in references [42, 43]) explore the dynamical systems analysis of interacting Umami chaplygin gas and interacting holographic dark energy in framework of adiabatic particle creation mechanism which show different cosmic phases of the universe. As nothing much is known about the physical nature of DE, probe is still on to hunt for an appropriate candidate for DE.
@@
On the other hand, interacting DE with varying mass DM particles can also provide a proper way to alleviate coincidence problem. The model of varying mass DM particles is based on the mechanism that decaying of DM reproduces the scalar field as DE. Here, mass of the dark matter is dependent on cosmic time ‘$t$’ via a scalar field ‘$\phi$’. It should be mentioned that this mechanism can be treated equivalent to an interaction included in the dark sectors, see for example [44, 45]. However, the model of varying mass has some basic differences from interacting models. The mass of dark matter can be exponential or power-law functions of scalar field or any other general form. Interacting quintessence studied recently with varying mass dark matter in the Lyra’s manifold in context of dynamical analysis, where mass of dark matter is assumed to be vary a exponential function of scalar field [46]. Interacting phantom is also studied where exponential and power-law of scalar field as well as exponential or power-law mass dependence is undertaken in the perspective of dynamical system [47]. A non-interacting scalar field model is investigated with arbitrary potential in Ref. [48] Consequently, a center manifold theory is employed to study the interacting phantom in association with varying mass dark matter [49]. A plethora of interacting quintessence models have been put forward over the years in the literature [50, 51, 52] where the dark matter mass depends on exponential or power-law of scalar field $\phi$, and these provided the possible solutions to coincidence problem. In this context, we would like to mention that the tachyon field can be made as a satisfactory candidate for the high energy inflation [16] and at the same time as an origin of DE depending on the form of the tachyon potential [20]. The DBI scalar field has been studied in context of modified symmetric teleparallel gravity in Ref. [53]. Also, some recent literature that studies cosmological application can be explored (for details see [54, 55]).
@@
The paper is organized as follows. In the next section II, we first discuss about the cosmological model considered here and then discuss about the formation of autonomous system and the relevant parameters. Phase space analysis has also been presented in section III. In addition, in section IV, we examine the behavior of diagnostic pair parameters. In section V we depicts the evolution of different cosmological quantities for our model and compare it with the observational data. Finally, our main findings and conclusions are shortly discussed in section VI. Throughout the paper, we use natural units in which $\kappa^{2}=8\pi G=c=1$.
@@ II.1 The model of interacting varying mass tachyonic DE
According to cosmological principle and from the survey of superclusters of galaxies [56], it is evident that the universe’s spatial section is homogeneous and isotropic at sufficiently large scales. Observation also predicts that the universe is spatially flat [78]. Thus, we consider the background universe which is in good agreement with observations, described by maximally symmetric space-time that is well described by the following flat Friedmann-Lemaître-Robertson-Walker (FLRW) metric [57]:
@@
where $\rho_{m}$ denotes the energy density of the DM. The energy density and the thermodynamic pressure for the tachyonic DE are denoted by $\rho_{\phi}$ and $p_{\phi}$ and are defined as follows [16, 17, 18, 19]:
@@
$Q=\frac{Q_{m}^{{}^{\prime}}(\phi)}{Q_{m}(\phi)}\dot{\phi}\rho_{m}=\tilde{\beta}(\phi)\dot{\phi}\rho_{m}$ (17)
where $\tilde{\beta}(\phi)=\frac{\partial}{\partial\phi}ln(Q_{m}(\phi))$. It should be noted that Honorez et al. [45] studied a coupled quintessence model in which the interaction with the dark matter sector is a function of the quintessence potential. They have also showed that such type of interaction (17) can arise from a field dependent mass term for the DM field. This kind of coupling (i.e., $\tilde{\beta}(\phi)=$ constant) can be found in the literature where non-minimally coupled Brans-Dicke theory containing a self-interacting potential, can automatically give rise to an interaction between the Brans-Dicke scalar field and the normal matter by applying a suitable conformal transformation (see [40] and the references therein).
Note that due to an interaction between the pressure-less DM and tachyon scalar field as DE, the evolution of the pressure-less DM does not follow the standard evolution $a^{-3}$. This is expected since an interaction will affect the evolution of the dark fluids. However, significant deviation from $\Lambda$CDM is not expected because if the coupling parameter is found to be large that may significantly affect the CMB spectrum. As $\Lambda$CDM is quite successful but still its revision is necessary according to its limitations from both observational and theoretical directions [58], therefore, one may wonder that the interacting scenario could be an approximate version of the $\Lambda$CDM model. Moreover, observational data also indicate that the coupling parameter is small and hence, interacting scenarios can explain the evolution of the universe at the level of perturbations [59, 60].
In our case, it is also clearly seen that the tachyonic scalar field and the DM do not evolve independently but interact with each other via an interaction term $Q$ in the energy conservation eqs. (15) and (16). It is important to note here that if we concentrate on the other interacting DE-DM models, the form of coupling chosen is ad-hoc and the source of such coupling is not known (see e.g. [36]). But, in our case, unlike other models, the interaction term is not an input but obtained its form from the field equations by considering the variable mass particle scenarios as already referred earlier. This makes our work very interesting and deserves further study in the present context.
Moreover, one can also express the conservation eqs. (15) and (16) in non-interacting form with effective EoS parameters for DM and DE as follows:
@@
The dynamics of cosmological evolution of the universe in this scenario of VAMP, we will perform the critical points analysis in the next section by transforming the cosmological evolution equations into an autonomous system of first order ordinary differential equations [32]:
$^{\prime}=\textbf{f(x)}$(x) (21)
[…]
$\xi^{\prime}=J\xi,$ (22)
where $J$ is the Jacobian matrix for the transformation, containing the coefficients of perturbation equations. The eigenvalues of the linearised Jacobian matrix will determine the nature of critical point $x=x_{c}$.
@@ II.2.1 Autonomous system with exponential mass function and exponential potential
In the model of exponential variable mass particles of DM, we consider the mass of DM is function of scalar field $\phi$ as $Q_{m}(\phi)=Q_{m0}~{}\mbox{exp}\{-\alpha H_{0}\phi\}$ and the potential of the scalar field $\phi$ as $V(\phi)=V_{0}~{}exp\{\beta H_{0}\phi\}$, where $Q_{m0}$ and $V_{0}$ are constant and $\alpha$ and $\beta$ are constant parameters. Now with the exponential potential and exponential mass function of DM, the 5D autonomous system (LABEL:autonomous_system_1) reduces to the following 3D autonomous system of ordinary differential equations:
@@
It is clear that the autonomous system (LABEL:autonomous_system) has singularities at $y=0$ and $z=1$. In this situation we can not study the properties of solutions lying on $y=0$ and $z=1$ plane. In order to remove the singularities, we multiply the right hand sides of the system (LABEL:autonomous_system) by $y^{2}(1-z)$. This operation allows us to analyze the solutions on $y=0$ and $z=1$ plane without changing the qualitative dynamical features of the system in the other regions of the phase space for $y^{2}(1-z)>0$. After applying this technique, we obtain
@@
It should be noted that the fig.1 is plotted for different values of parameters ($\alpha$ and $\beta$) with different initial conditions, and it has been shown that the model does not exhibit any negative energy density during the evolution of the universe for a wide range of coupling. Thus, the present interacting tachyon model behaves well in cosmological point in view, where the interaction $Q$ ensures physically consistent energy transfer, such as consistent with conservation laws, avoiding negative energy density and instabilities in cosmic evolution.
Figure 1: The figures show the evolution of density parameters for tachyon and dark matter for different initial conditions. Panel (a) for $\alpha=0.5,~{}\beta=0.5$. Panel (b) for $\alpha=0.001,~{}\beta=0.1$. Panel (c) for $\alpha=-0.001,~{}\beta=-0.1$.
The effective equation of state parameter of tachyon field (DE) for this model read as
@@
From the above, one can achieve the condition for acceleration of the universe when: $q<0$, i.e., when $\omega_{eff}<-\frac{1}{3}$ and for deceleration, one follows the condition: $q>0$, i.e., $\omega_{eff}>-\frac{1}{3}.$ Friedmann eq. (2) will give the constraint equation of $\Omega_{m}$ explicitly depending on the variables $x$ and $y$ as
@@
III Phase space analysis of autonomous system (LABEL:autonomous_system_2):
The critical points for the system (LABEL:autonomous_system_2) are the following
@@ III Phase space analysis of autonomous system (LABEL:autonomous_system_2):
For autonomous system (LABEL:autonomous_system_2), Critical points and their corresponding physical parameters are shown in the table 1.
Table 1: The Critical points and the corresponding physical parameters for the system (LABEL:autonomous_system_2) are presented
@@
Table 2: The eigenvalues of the linearized system (LABEL:autonomous_system_2)
@@
The table 2 shows the eigenvalues of the Jacobian matrix (evaluated at the critical points) for the autonomous system (LABEL:autonomous_system_2) and we note that all the critical critical points are non-hyperbolic in nature. So we can not analyze them by linear stability theory. We need to use center manifold theory for some cases. But center manifold theory can be used only for the critical points $B$ and $C$ as rest of the critical points do not have any eigenvalues with nonzero real part. To analyze $A,D,E,F,G$ we use numerical procedure.
@@
It is to be noted that the origin is a saddle on the $xy$ plane (see figure 2) and the vector field near $x$-axis ($A$) is unstable on both $xy$ and $xz$ plane (see figure 2). The vector field near the critical point $D$ on the $xy$-plane behaves as unstable node. The flow is repelling along $x$-axis and $y$-axis on the $xy$ plane. On the other hand, the line $(1,0,z_{c})$ is an attractor for $z_{c}>0$ and repeller for $z_{c}<0$ on the $xz$ plane. But the line $(1,0,z_{c})$ is a repeller on the $yz$ plane. In figure (4) it is to be noted that the vector field on the $xy$ plane near the critical point $E$ is unstable in nature. On the $xz$ plane the vector field is attracting for $z_{c}<0$ and repelling for $z_{c}>0$.
@@
$\bullet$ The set of critical points $B$ and $C$ exist for all $\beta\in\mathbb{R}$ and $\alpha$. These solutions are completely dominated by tachyon scalar field $(\Omega_{\phi}=1)$ which behaves as cosmological constant ($\omega^{\phi}_{eff}=-1$). The cosmic evolution near the points characterizes the de-Sitter expansion ($q=-1$) of the universe. These critical points are isolated non-hyperbolic in nature and the Jacobian matrix at $B$ or $C$ contains only one vanishing eigenvalue. We can employ center manifold theory (see for instance, [61, 62, 63, 64] ) to find the characteristic of the vector field near these critical points as the following:
@@
The equation (LABEL:autonomous_system_2) reduces to
$\frac{dx}{dN}$ $=\left(1-x^{2}\right)\left(\alpha\sqrt{1-x^{2}}z+3x(y+1)^{2}(z-1)-(y+1)^{2}z(\alpha+\beta)\right).$ (35a)
$\frac{dy}{dN}$ $=\frac{1}{2}(y+1)^{3}\left(3(1-z)\left(1-\sqrt{1-x^{2}}(y+1)^{2}\right)+\beta xz\right).$ (35b)
$\frac{dz}{dN}$ $=\frac{3}{2}(y+1)^{2}(z-1)^{2}z\left(1-\sqrt{1-x^{2}}(y+1)^{2}\right).$ (35c)
The Jacobian matrix of the system (35a-35c) can be obtained as
$J(x,y,z)=\begin{bmatrix}xzf-9x^{2}h(z-1)+3h(z-1)&2\left(1-x^{2}\right)(y+1)(3x(z-1)-z(\alpha+\beta))&\left(1-x^{2}\right)\left(\alpha g+3xh-h(\alpha+\beta)\right)\\ \frac{1}{2}(y+1)^{3}\left(\beta z-\frac{3xh(z-1)}{g}\right)&\frac{3}{2}h\left(2gh(z-1)+3(1-z)\left(1-gh\right)+\beta xz\right)&\frac{1}{2}(y+1)^{3}\left(\beta x-3\left(1-gh\right)\right)\\ \frac{3x(y+1)^{4}(z-1)^{2}z}{2g}&3(y+1)(z-1)^{2}z\left(1-2gh\right)&-\frac{3}{2}h(z-1)(3z-1)\left(gy(y+2)+g-1\right)\\ \end{bmatrix}$
where, $f=\left(2h(\alpha+\beta)-3\alpha g\right)$, $g=\sqrt{1-x^{2}}$, $h=(y+1)^{2}$ and at the origin the matrix takes the following simplified form
$J(0,0,0)=\begin{bmatrix}-3&0&-\beta\\ 0&-3&0\\ 0&0&0\end{bmatrix}.$
To find center manifold at the origin, we need to diagonalize the above Jacobian matrix. For that we need to find $T$ such that $T^{-1}JT$ is a diagonal matrix. So $T$ is given by
$T=\begin{bmatrix}1&0&-\frac{\beta}{3}\\ 0&1&0\\ 0&0&1\end{bmatrix}$
If we consider J(0,0,0) as a linear transformation from $\mathbb{R}^{3}$ to $\mathbb{R}^{3}$ and the basis of $\mathbb{R}^{3}$ changes by $T$ matrix, then the coordinate vector (variables) $(x,y,z)$ transform to $(P,Q,R)$-coordinate system by the following rule
$\begin{bmatrix}x\\ y\\ z\end{bmatrix}=\begin{bmatrix}P-\frac{\beta}{3}R\\ Q\\ R\end{bmatrix}$ (36)
@@
$\begin{bmatrix}P\\ Q\\ R\end{bmatrix}=\begin{bmatrix}x+\frac{\beta}{3}z\\ y\\ z\end{bmatrix}$ (37)
Now in the new $(P,Q,R)$-coordinate system the autonomous equations (35a-35c) take the following form:
$\frac{dP}{dN}=\frac{1}{2}\beta(Q+1)^{2}(R-1)^{2}R\left(1-(Q+1)^{2}\sqrt{1-\left(P-\frac{\beta R}{3}\right)^{2}}\right)\\ -\frac{1}{27}\left((\beta R-3P)^{2}-9\right)\left(9P(Q+1)^{2}(R-1)+\alpha R\sqrt{9-(\beta R-3P)^{2}}-3(Q+1)^{2}R(\alpha+\beta R)\right).$ (38)
$\frac{dQ}{dN}$ $=\frac{1}{2}(Q+1)^{3}\left(3(1-R)\left(1-(Q+1)^{2}\sqrt{1-\left(P-\frac{\beta R}{3}\right)^{2}}\right)+\beta R\left(P-\frac{\beta R}{3}\right)\right).$ (39)
$\frac{dR}{dN}$ $=\frac{3}{2}(Q+1)^{2}(R-1)^{2}R\left(1-(Q+1)^{2}\sqrt{1-\left(P-\frac{\beta R}{3}\right)^{2}}\right).$ (40)
Now near the origin higher degree terms of the equations of the vector field do not contribute significantly. So ignoring the higher degree terms in the equations (38-40) one can obtain following simplified form
$\frac{dP}{dN}$ $=-3P+3PR-6PQ-(2\alpha+\beta)QR+\text{~{}higher degree terms}.$ (41)
$\frac{dQ}{dN}$ $=-3Q-\frac{21}{2}Q^{2}+3QR+\frac{\beta}{2}PR+\text{~{}higher degree terms}.$ (42)
$\frac{dR}{dN}$ $=6R^{2}Q-\frac{15}{2}Q^{2}R+\text{~{}higher degree terms}.$ (43)
Then the flow along the center manifold at $B$ is topologically equivalent (as $\det(T)>0$) to the flow of
$\frac{dR}{dN}\approx 0.$ (44)